The Hamilton Corporation has 5 million shares of stock outstanding and will report earnings of $6,460,000 in the current year. The company is considering the issuance of 2 million additional shares which can only be issued at $35 per share.


a. Assume the Hamilton Corporation can earn 8.00 percent on the proceeds. Calculate the earnings per share. (Do not round intermediate calculations and round your answer to 2 decimal places.)

Earnings per Share -


b. Should the new issue be undertaken based on earnings per share?
  

Yes
No
Answer #1

Hello,
a) EPS = Earning per share = Earningsavailable forequity Noofequity shares

Earnings = $6460000
Outstanding shares = 5 million
Earnings on new shares issue = 2000000*35 = 70000000 * 8%= 5600000
Total Earnings = 6460000 + 5600000= 12060000
Total outstanding shares = 5 + 2 million= 7 million

EPS = 12060000 / 7000000 = 1.72286


b) EPS before additional share issued = 6460000 / 5000000 = 1.29
EPS in case of additional share issued = 5600000 / 2000000 = 2.8
EPS is increasing while additional share issued

Answer of Part (b) is yes

I hope this clear your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.